OpenStudy (anonymous):

Could someone help me on this question please: If g(x) = 3+x+e^x, find g^-1(4) Answer: g^-1(4) = 0

OpenStudy (anonymous):

I replaced all x's with y's and came up with this, x-3=y+e^y and I'm not sure how to solve for y from here

OpenStudy (anonymous):

*all x's with y's and all y's with x's

OpenStudy (ash2326):

You need to find \[g ^{-1} (4)\] This means you need to find x for which g(x)=4 You agree with this?

OpenStudy (anonymous):

Yeah, that makes sense so far

OpenStudy (anonymous):

Though how would I do that? Is that just plug and check?

OpenStudy (ash2326):

so \[4=3+x+e^x\] Yeah here if we plug x=0 We'll get 4 so x=0 is the answer

OpenStudy (ash2326):

I think we can only plug and check:(

OpenStudy (anonymous):

Oh okay, thanks for your help, I appreciate it.